The monthly market demand curve for calculators among engineering students is given by 

P = 100 – Q. where P is the price per calculator in dollars and Q is the number of calculators purchased per month. If the price is UD$ 30, how much revenue will calculator makers get each month? Find the price elasticity of demand for calculators at this point. What should calculator makers do to increase revenue?



1
Expert's answer
2021-06-17T09:44:44-0400

"P=100-Q\\Q=100-P"


substituting P=UD$ 30


"Q=100-30\\Q=70"


"Ped=\frac{\Delta Q}{\Delta P}\times \frac{P}{Q}"


"\frac{\Delta Q}{\Delta P}=-1"


"Ped=-1\times\frac{30}{70}=-0.429"


Ped<1

The demand is inelastic and therefore a change in price will have little or no effect on the demanded quantity hence to increase revenue the price should be increased